OpenStudy (anonymous):

Please assist me? I'll answer anything related to English, Psychology, or Philosophy in return? ;) What is the equation of the line passing through (2, -3) and parallel to the line y = -6x - 1 in standard form? a.) 6x + y = 9 b.) 6x - y = -9 c.) x + 6y = 9 d.) x - 6y = -9

OpenStudy (anonymous):

parallel to \(y=-6x-1\) means slope will be \(-6\) as well we can start with \[y+3=-6(x-2)\] multiply out and ge t \[y+3=-6x+12\] or \[6x+y=9\]

OpenStudy (anonymous):

How'd you get 12?

OpenStudy (anonymous):

distributive law

OpenStudy (anonymous):

\[-6(x+2)=-6\times x+(-6)\times(- 2)=-6x+12\]

OpenStudy (anonymous):

So, like an examples would be... 2 × (1 + 3) = (2 × 1) + (2 × 3) but 2 /(1 + 3) is not equal to (2 / 1) + (2 / 3).

OpenStudy (anonymous):

an example*

OpenStudy (anonymous):

typo \[-6(x-2)=-6\times x+(-6)\times(- 2)=-6x+12\]

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

yes your example is right yw